OpenStudy (anonymous):

plz help asap If A = {2, 4, 6, 8, 10} and B = {1, 2, 3, 4, 5}, find A Intersects with B. A. {2, 4} B. {1, 2, 3, 4, 5, 6, 8, 10} C. {7, 9} D. {1, 3, 5}

OpenStudy (freckles):

just identity the elements that are in BOTH A and B (what do they have in common; what elements)

OpenStudy (anonymous):

i dont understand it though i thought it was a but got it wrong

OpenStudy (freckles):

you thought it was what?

OpenStudy (anonymous):

A

OpenStudy (freckles):

that is right both A and B have 2 and 4 in common

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

i must have clicked on the wrong one

OpenStudy (freckles):

\[\cap \text{ is intersect } \\ \cup \text{ is union }\]

OpenStudy (freckles):

you did see the first symbol right?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

can you help me some more

OpenStudy (freckles):

sure

OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. A. x + 6y = 8 y = 2x – 8 B. y = 2x – 7 x + 2y = 3 C. 2y = –3x + 5 2x + 3y = 4 D. 5x – 8y = 9 12x – 5y = 7

OpenStudy (anonymous):

idk how to find if there perpindiculer

OpenStudy (anonymous):

hello?

geerky42 (geerky42):

Lines are perpendicular if \(m_1~m_2 = -1\) (product of slopes of these lines has to be -1.)

geerky42 (geerky42):

Let's take a look at A. What is slope of first equation? Second equation?

OpenStudy (anonymous):

idk i think 3